Bialya’s currency is the dinar (BD). Kahndaq’s currency is the riyal (KR). Please calculate the Bialya/Kahndaq exchange rate., EBD/KR. Bialya has a money supply of BD 80 trillion. Bialya’s gross domestic product (GDP) is BD 4 trillion. Bialya’s liquidity constant is 2. Kahndaq’s money supply is KR 10 trillion. Kahndaq’s gross domestic product is K 2 trillion. Kahndaq’s liquidity constant is also 2. Be sure to include the relevant equation and calculations in your answer. Hint 1: You will know which equation to use from the information provided. Hint 2: GDP = Y Hint 3: Since every number is in the trillions, you just need to calculate the numbers given, without worrying about the trillions part
Answer #1

Answer

Bialya’s currency is the dinar (BD). Kahndaq’s currency is the riyal (KR). Please calculate the Bialya/Kahndaq exchange rate., EBD/KR. Bialya has a money supply of BD 80 trillion.

First we will calculate the price level in two countries using the quantity theory of money. And then, we will use law of one price to determine exchange rate.

The quantity theory is money says that: M*V= P*Y

Where M is money supply, V is velocity of money /liquidity constant, P is price level and Y is GDP.

In Bialya, Mb*Vb= Pb*Yb => Pb = Mb*Vb/Yb = 80*2/4 = 160/4=40

In Kahndaq, Mk*Vk=Pk*Yk =>Pk= Mk*Vk/Yk =10*2/2= 10

Thus, price level in Bialya is 40 and in Kahndaq, its 10.

Now, according to the law of one price,

Nominal exchange rate is the ratio of prices in two countries.

Thus, EBD/KR= Pb/Pk = 40/10 = 4

Which means that 1BD= 4KR.

IF YOU HAVE ANY QUERY PLEASE COMMENT DOWN BELOW
PLEASE GIVE A THUMBS UP